Theory and Applications of Categories, Vol. 32, No. 41, 2017, pp. 1454–1484.

TOPOLOGICAL PROPERTIES OF NON-ARCHIMEDEAN APPROACH SPACES

EVA COLEBUNDERS AND KAREN VAN OPDENBOSCH

Abstract. In this paper we give an isomorphic description of the category of non-Archimedian approach spaces as a category of lax algebras for the ultrafilter monad and an appropriate quantale. Non-Archimedean approach spaces are characterised as those approach spaces having a tower consisting of topologies. We study topological properties p, for p compactness and Hausdorff separation along with low-separation properties, regularity, normality and extremal disconnectedness and link these properties to the condition that all or some of the level topologies in the tower have p. A compactification technique is developed based on Shanin’s method.

1. Introduction

In monoidal topology [Hofmann, Seal, Tholen (eds.), 2014] starting from the quantale P_{+} = ([0, ∞]^{op}, +, 0) where the natural order ≤ of the extended real halfline [0, ∞] is reversed so that 0 = ⊤ and ∞ = ⊥, based on Lawvere’s description of quasi-metric spaces as categories enriched over the extended real line [Lawvere 1973], one obtains an isomorphism between the category qMet of extended quasi-metric spaces with non-expansive maps and P_{+}-Cat, the category of P_{+}-spaces (X, a), where X is a set and a : X⟶ X is a transitive and reflexive P_{+}-relation, meaning it is a map a : X × X → P_{+} satisfying a(x, y) ≤ a(x, z) + a(z, y) and a(x, x) = 0 for all x, y, z in X, with morphisms f : (X, a) → (Y, b) of P_{+}-Cat, maps satisfying b(f(x), f(x′)) ≤ a(x, x′) for all x, x′ in X.

TOPOLOGICAL PROPERTIES OF NON-ARCHIMEDEAN APPROACH SPACES 1455

Our interest goes to the larger category App of approach spaces and contractions, which contains qMet as coreflectively embedded full subcategory and which is better behaved than qMet with respect to products of underlying topologies. A comprehensive source on approach spaces and their vast field of applications is [Lowen, 2015] in which several equivalent descriptions of approach spaces are formulated in terms of (among others) distances, limit operators, towers and gauges. A first lax-algebraic description of approach spaces was established by Clementino and Hofmann in [Clementino, Hofmann, 2003]. The construction involves the ultrafilter monad β = (β, m, e), the quantale P_{+} and an extension of the ultrafilter monad to P_{+}-relations. The so constructed category (β, P_{+})-Cat and its isomorphic description App became an important example in the development of monoidal topology [Hofmann, Seal, Tholen (eds.), 2014]. More details on the construction will be recalled in section 2.

Received by the editors 2017-07-17 and, in final form, 2017-11-01.

Transmitted by Walter Tholen. Published on 2017-11-13.

2010 Mathematics Subject Classification: 18C15, 18C20, 54A05, 54B30, 54E99.

Key words and phrases: Lax algebra, quantale, non-Archimedean approach space, quasi-ultrametric space, initially dense object, topological properties in (β, P_{∨})-Cat, compactification.

\begin{matrix}c \\ ◯ \end{matrix} Eva Colebunders and Karen Van Opdenbosch, 2017. Permission to copy for private use granted.

In this paper we consider the non-Archimedian counterparts of the previous well known constructions. These are obtained by switching the quantale to P_{∨}. As before we consider the extended real halfline [0, ∞] as a complete lattice with respect to the natural order ≤ and we reverse its order, so that 0 = ⊤ is the top and ∞ = ⊥ is the bottom element. Since [0, ∞]^{op} is a chain, it is a frame, and we may consider it a quantale with its meet operation which, according to our conventions on the reversed order, is the supremum with respect to the natural order of [0, ∞] and as in approach theory this operation will be denoted by ∨ (note that here we deviate from the notation in [Hofmann, Seal, Tholen (eds.), 2014]). As stated in exercise III.2.B of [Hofmann, Seal, Tholen (eds.), 2014] the category P_{∨}-Cat is isomorphic to the category qMet^{u} of extended quasi-ultrametric spaces (X, d) with nonexpansive maps, where an extended quasi-metric is called an extended quasi-ultrametric if it satisfies the strong triangular inequality d(x, z) ≤ d(x, y) ∨ d(y, z), for all x, y, z ∈ X.

In section 2, using the ultrafilter monad β = (β, m, e), with P_{∨} as quantale and an extension of the ultrafilter monad to P_{∨}-relations we show that the category of associated lax algebras (β, P_{∨})-Cat is isomorphic to the reflectively embedded full subcategory NA-App of App, the objects of which are called non-Archimedean approach spaces. In section 3 we give equivalent descriptions of non-Archimedean approach spaces in terms of distances, limit operators, towers and gauges. The characterization in terms of towers can be easily expressed: an approach space is non-Archimedean if and only if its tower (T_{ε})_{ε∈ℝ^{+}} consists of topologies. We show that NA-App can be generated as the concretely reflective hull by one particular extended quasi-ultrametric.

Following the general philosophy to consider lax algebras as spaces, it is our main purpose in this paper to study topological properties p like compactness and Hausdorff separation, along with low-separation properties, regularity, normality and extremal disconnectedness for non-Archimedean approach spaces. In section 6 we introduce these properties (β, P_{∨})-p as an application to (β, P_{∨})-Cat of the corresponding (𝕋, V)-properties for arbitrary (𝕋, V)-Cat as developed in V.1 and V.2 of [Hofmann, Seal, Tholen (eds.), 2014]. For each of the properties p we characterise (β, P_{∨})-p in the context of NA-App. On the other hand we make use of the well known meaning of these properties in the setting of Top and for a non-Archimedean approach space X with tower of topologies (T_{ε})_{ε∈ℝ^{+}} we compare the property (β, P_{∨})-p to the properties X has p at level 0, meaning the topological space (X, T_{0}) has p in Top, X almost strongly has p, meaning (X, T_{ε}) has p in Top for every ε ∈ ℝ_{0}^{+} and X strongly has p, meaning (X, T_{ε}) has p in Top for every ε ∈ ℝ^{+}. For p the Hausdorff property, low separation properties or regularity the conditions strongly p, almost strongly p and (β, P_{∨})-p are all equivalent and different from p at level 0. For p compactness strongly p is equivalent to p at level 0, and almost strongly p is equivalent to (β, P_{∨})-p. For p normality or extremal disconnectedness strongly p implies almost strongly p which implies (β, P_{∨})-p. Moreover strongly p implies p at level 0. We give counterexamples showing that there are no other valid implications.

In the last sections of the paper we develop a compactification technique and prove that every non-Archimedean approach space X given by its tower of topologies (T_{ε})_{ε∈ℝ^{+}} can be densely embedded in a compact (at level 0) non-Archimedean approach space Y given by its tower of topologies (S_{ε})_{ε∈ℝ^{+}} in such a way that at level 0 the compact space (Y, S_{0}) is the Shanin compactification of (X, T_{0}). We give necessary and sufficient conditions for the compactification to be Hausdorff at level 0 and we prove that a space X that is Hausdorff at level 0, (β, P_{∨})-regular and (β, P_{∨})-normal has such a Hausdorff compactification.

2. P_{∨}-Cat and (β, P_{∨})-Cat

We follow definitions and notations from [Hofmann, Seal, Tholen (eds.), 2014], unless stated otherwise. In this section we adapt the constructions of P_{+}-Cat and (β, P_{+})-Cat in [Hofmann, Seal, Tholen (eds.), 2014] to the quantale P_{∨}. Consider the extended real halfline [0, ∞] which is a complete lattice with respect to the natural order ≤. We reverse its order, so that 0 = ⊤ is the top and ∞ = ⊥ is the bottom element. When working with ([0, ∞]^{op}, ≤^{op}) and forming infima or suprema we will denote these by inf^{op}, ⋀^{op}, sup^{op} or ⋁^{op}. It means that we will deviate from the conventions made in [Hofmann, Seal, Tholen (eds.), 2014] since we will use both the symbols inf and ⋀ when forming infima and sup and ⋁ when forming suprema, referring to the natural order on ([0, ∞], ≤).

Since [0, ∞]^{op} is a chain, it is a frame, and we may consider it a quantale P_{∨} = ([0, ∞]^{op}, ∨, 0) with its meet operation (which, according to our conventions, is the supremum with respect to the natural order of [0, ∞] and will be denoted by ∨). The map a ∨ (−) : P_{∨} ⟶ P_{∨} is left adjoint to the map a−•(−) : P_{∨} ⟶ P_{∨} defined by

a−•b = inf\{v ∈ [0, ∞] | b ≤ a ∨ v\} = \{ \begin{matrix}0 & a ≥ b, \\ b & a < b. \end{matrix}

The category P_{∨}-Rel has sets as objects and P_{∨}-relations as morphisms. A P_{∨}-relation r : X⟶ Y from X to Y is represented by a map r : X × Y → P_{∨}. Two P_{∨}-relations r : X⟶ Y and s : Y ⟶ Z can be composed in the following way

(s · r)(x, z) = inf {r(x, y) ∨ s(y, z) | y ∈ Y }.

2.1. Definition. P_{∨}-Cat is the category of P_{∨}-spaces (X, a), where X is a set and a : X⟶ X is transitive and reflexive P_{∨}-relation, meaning it is a map a : X × X → P_{∨} satisfying

a(x, y) ≤ a(x, z) ∨ a(z, y) and a(x, x) = 0

for all x, y, z in X. A morphism f : (X, a) → (Y, b) of P_{∨}-Cat is a map satisfying b(f(x_{1}), f(x_{2})) ≤ a(x_{1}, x_{2}) for all x_{1}, x_{2} in X.

The map φ : P_{∨} ⟶ P_{+} with φ(v) = v, for all v ∈ P_{∨}, is a lax homomorphisms of quantales. This induces a lax-functor φ : P_{∨}-Rel ⟶ P_{+}-Rel, which leaves objects unaltered and sends r : X × Y ⟶ P_{∨} to φ · r : X × Y ⟶ P_{+}. Obviously φ is compatible with the identical lax extension of the identity monad 𝕀 to P_{∨}-Rel and P_{+}-Rel. Hence, this lax-functor induces a change-of-base functor B_{φ} : P_{∨}-Cat ⟶ P_{+}-Cat. Moreover, this change-of-base functor is an embedding. Whereas P_{+}-Cat ≅ qMet, the category of extended quasi-metric spaces and non-expansive maps, it is known that P_{∨}-Cat ≅ qMet^{u}, the category of quasiultrametric spaces. An early systematic study of ultrametric spaces can be found in the work of Monna [Monna] and of de Groot [de Groot] in the 50’s, but ever since the amount of literature has become extensive as (quasi)-ultrametrics became important tools in a wide range of fields, like combinatorics, domain theory, linguistics, solid state physics, taxonomy and evolutionary tree constructions, to name only a few.

For a set X let βX be the set of all ultrafilters on X and for a map f : X ⟶ Y, let βf : βX ⟶ βY : U ↦ {B ⊆ Y | f^{−1}(B) ∈ U}. Then β : Set ⟶ Set is a functor. To avoid overloaded notations, we will often write f(U) instead of βf(U). Now consider the ultrafilter monad β = (β, m, e) on Set with components m_{X} : ββX → βX defined by the Kowalsky sum of 𝔛 ∈ ββX

m_{X}𝔛 = Σ𝔛 = {A ⊆ X | {U ∈ βX | A ∈ U} ∈ 𝔛},

and e_{X} : X ⟶ βX defined by e_{X}(x) = ẋ = {A ⊆ X | x ∈ A} for x ∈ X.

Analogous to the P_{+} situation, for a P_{∨}-relation r : X⟶ Y and α ∈ [0, ∞], we can define the relation r_{α} : X⟶ Y by x r_{α} y ⇔ r(x, y) ≤ α. For A ⊆ X we put r_{α}(A) = {y ∈ Y | ∃x ∈ A : x r_{α} y}, and for A ⊆ P(X) we let r_{α}(A) = {r_{α}(A) | A ∈ A}. Recall that β : Set → Set can be extended to a 2-functor \overline{β} : Rel → Rel where for any sets X and Y, any relation r : X⟶ Y and ultrafilters U ∈ βX and W ∈ βY we have

U (\overline{β}r) W ⇔ r(U) ⊆ W.

Then, for r : X⟶ Y a P_{∨}-relation, we let

\overline{β}r(U, W) = inf\{α ∈ [0, ∞] | U\overline{β}(r_{α})W\},

for U ∈ βX and W ∈ βY. Analogous to the P_{+} situation, the extension of the ultrafilter monad β is a flat and associative lax extension to P_{∨}-Rel, which we again denote by \overline{β} = (\overline{β}, m, e).

2.2. Definition. Let (β, P_{∨})-Cat be the category of lax algebras (X, a) for the P_{∨}-Rel-extension of the ultrafilter monad, where X is a set and a : βX⟶ X is a P_{∨}-relation that is a map a : βX × X → P_{∨} that is reflexive, meaning

a(ẋ, x) = 0

for all x ∈ X and transitive, meaning

a(m_{X}(𝔛), x) ≤ \overline{β}a(𝔛, U) ∨ a(U, x)

for all 𝔛 ∈ ββX, for all U ∈ βX and for all x ∈ X. A morphism f : (X, a) → (Y, b) of (β, P_{∨})-Cat is a map satisfying

b(f(U), f(x)) ≤ a(U, x)

for all x in X and U ∈ βX.

It is clear that the lax-homomorphism of quantales φ : P_{∨} ⟶ P_{+} is compatible with the lax-extension of the ultrafilter monad β to P_{∨}-Rel and P_{+}-Rel. Hence, this induces another change-of-base functor

C_{φ} : (β, P_{∨})-Cat ⟶ (β, P_{+})-Cat.

Again, this change-of-base functor is an embedding. The category (β, P_{+})-Cat is known to be isomorphic to App . This result was first shown in [Clementino, Hofmann, 2003] and can be found in section III.2.4 in [Hofmann, Seal, Tholen (eds.), 2014]. Both proofs go via distances. We will proceed by giving an isomorphic description of the category (β, P_{∨})-Cat and in order to do so we introduce the concept of non-Archimedean approach spaces by strengthening axiom (LU*) in [Lowen, 2015].

2.3. Definition. Let NA-App be the full subcategory of App consisting of approach spaces (X, λ), with a limit operator λ : βX → PX_{∨} satisfying the strong inequality

(Lβ_{∨}^{∗}) For any set J, for any ψ : J ⟶ X, for any σ : J ⟶ βX and for any U ∈ βJ

λΣσ(U) ≤ λψ(U) ∨ sup_{U∈U} inf_{j∈U} λσ(j)(ψ(j)).

We will call these objects non-Archimedean approach spaces and the structure λ will be called a non-Archimedean limit operator.

Most of the time we will be working with ultrafilters. However we should mention that one can equivalently define NA-App as a full subcategory of App by strengthening axiom (L*) in [Lowen, 2015] for a limit operator on the set FX of all filters on X, again by replacing the + on the righthand side of the inequality by ∨. In terms of filters the condition can be proven to be equivalent to property [F_{s}] in [Brock, Kent, 1998].

2.4. Theorem. The category (β, P_{∨})-Cat of lax algebras for the P_{∨}-Rel-extension of the ultrafilter monad is isomorphic to NA-App.

Proof. The isomorphism directly links an approach space (X, λ) with λ : βX → PX_{∨} to (X, a) where a is the corresponding map a : βX × X → P_{∨} and vice versa. As NA-App is a subcategory of App, the limit operator λ satisfies λ(ẋ)(x) = 0, for all x ∈ X. So this corresponds to reflexivity of a. In order to show that the axiom (Lβ_{∨}^{∗}) for λ corresponds to transitivity of a, we can use similar techniques as in the proof of theorem 12.7, (β, P_{+})-Cat ≅ App, which can be found in [Lowen, 2015]. Finally that through the identification of lax algebraic structures and non-Archimedean limit operators, also morphisms in both categories coincide, follows from the characterization of contractions in App via ultrafilters.

3. Equivalent descriptions of non-Archimedean approach spaces

In this section we define various equivalent characterizations for non-Archimedean approach spaces, in terms of distances, towers and gauges.

3.1. Non-Archimedean distances. First we associate to a limit operator λ of an approach space X, its unambiguously defined distance δ : X × 2^{X} ⟶ [0, ∞]. This operator provides us with a notion of distance between points and sets. The smaller the value of δ(x, A) the closer the point x is to the set A. When starting with a non-Archimedean limit operator, we get a non-Archimedean distance, i.e. a distance which satisfies a stronger triangular inequality. Since the proof of the next proposition is a modification of the proof in [Lowen, 2015], replacing + by ∨ and since an analogous result for [F_{s}] appears in [Brock, Kent, 1998], we omit it here.

3.2. Proposition. If λ : βX ⟶ PX_{∨} is the limit operator of a non-Archimedean approach space X, then the associated distance, given by

δ : X × 2^{X} ⟶ P_{∨} : (x, A) ↦ inf_{U∈βA} λU(x),

satisfies the strong triangular inequality (D4_{∨})

δ(x, A) ≤ δ(x, A^{(ε)}) ∨ ϵ,

for all x ∈ X, A ⊆ X and ε ∈ P_{∨}, with A^{(ε)} := {x ∈ X | δ(x, A) ≤ ϵ}.

If δ : X × 2^{X} ⟶ P_{∨} is a distance of an approach space X satisfying the strong triangular inequality (D4_{∨}), then the associated limit operator, given by

λ : βX ⟶ PX_{∨} : U ↦ sup_{U∈U} δ_{U},

is a non-Archimedean limit operator.

Distances satisfying the strong triangular inequality (D4_{∨}) are called non-Archimedean distances. The following inequality will be useful later on and has a straightforward proof.

3.3. Proposition. If δ : X × 2^{X} ⟶ P_{∨} is a non-Archimedean distance, then the following inequality holds.

δ(x, A) ≤ δ(x, B) ∨ sup_{b∈B} δ(b, A),

for all x ∈ X and A, B ⊆ X.

We give an example of a non-Archimedean approach space on P_{∨} which will play an important role later on in Section 5.

3.4. Example. Define δ_{P_{∨}} : P_{∨} × 2^{P_{∨}} ⟶ P_{∨} by

δ_{P_{∨}}(x, A) := \{ \begin{matrix}sup A−•x & A ≠ ∅, \\ ∞ & A = ∅. \end{matrix}

= \{ \begin{matrix}0 & A ≠ ∅ and x ≤ sup A, \\ x & A ≠ ∅ and x > sup A, \\ ∞ & A = ∅. \end{matrix}

Then δ_{P_{∨}} is a non-Archimedean distance on P_{∨}. The associated non-Archimedean limit operator is determined as follows. Given an ultrafilter U on P_{∨} and an element x ∈ X, then

λ_{P_{∨}}U(x) = sup_{U∈U}(sup U−•x)

= \{ \begin{matrix}0 & ∀U ∈ U : x ≤ sup U, \\ x & ∃U ∈ U : x > sup U. \end{matrix}

3.5. Non-Archimedean towers. To a distance of an approach space X, we now associate its unambiguously defined tower (𝔱_{ε})_{ε∈ℝ^{+}} of closure operators. Linking the limit operator on filters satisfying [F_{s}] to the tower, an analogous result appears in [Brock, Kent, 1998], so here we omit the proof.

3.6. Proposition. If δ : X × 2^{X} ⟶ P_{∨} is the distance of a non-Archimedean approach space X, then all levels of the associated tower (𝔱_{ε})_{ε∈ℝ^{+}} , with

𝔱_{ε}(A) = A^{(ε)}, ∀A ⊂ X, ∀ε ∈ ℝ^{+},

are topological closure operators.

If (𝔱_{ε})_{ε∈ℝ^{+}} is the tower of an approach space X, where all levels are topological closure operators, then the associated distance

δ : X × 2^{X} ⟶ P_{∨} : (x, A) ↦ inf{ε ∈ ℝ^{+} | x ∈ 𝔱_{ε}(A)}

is a non-Archimedean distance on X.

Towers of approach spaces that satisfy the stronger condition of the previous proposition are called non-Archimedean towers. At each level the closure operator 𝔱_{ε} defines a topology T_{ε}, so we will denote the structure also by (T_{ε})_{ε∈ℝ^{+}} , when working with open sets at each level, or (C_{ε})_{ε∈ℝ^{+}} , when using closed sets.

From the characterisation of approach spaces in terms of towers given in [Lowen, 2015] we now have the following.

3.7. Corollary. A collection (T_{ε})_{ε∈ℝ^{+}} of topologies on a set X defines a tower for some non-Archimedean approach space if and only if it satisfies the coherence condition

T_{ε} = ⋁_{γ>ε} T_{γ},

where the supremum is taken in Top.

We include more examples of non-Archimedean approach spaces that will be useful in section 6. The construction of the non-Archimedean approach spaces in the following examples is based on 3.7.

3.8. Example. Let X be a set and S a given topology on X. Let (T_{ε})_{ε∈ℝ^{+}} be defined by

T_{ε} = \{ \begin{matrix}P(X) whenever 0 ≤ ε < 1 \\ S  whenever 1 ≤ ε < 2 \\ \{X, ∅\} whenever 2 ≤ ε \end{matrix}

Clearly the coherence condition is satisfied and so (X, (T_{ε})_{ε∈ℝ^{+}} ) defines a non-Archimedean approach space which we will denote by X_{S}.

3.9. Example. Let X =]0, ∞[, endowed with a topology T with neighborhood filters (V(x))_{x∈X} and assume T is finer than the right order topology. We define (T_{ε})_{ε∈ℝ^{+}} with T_{0} = T and T_{ε} at level 0 < ε having a neighborhood filter

V_{ε}(x) = \{ \begin{matrix}\{X\} whenever x ≤ ε \\ V(x) whenever ε < x \end{matrix}

at x ∈ X. For a fixed level ε and ε < x the set {G ∈ T | x ∈ G, G ⊆]ε, ∞[} is an open base of V_{ε}(x). So clearly (T_{ε})_{ε∈ℝ^{+}} is a decending chain of topologies. To check the other inclusion of the coherence condition, let 0 ≤ ε and x ∈]0, ∞[. Either x ≤ ε and then V_{ε}(x) = {X} ⊆ V_{γ}(x) for every γ. Or ε < x, then choose γ with ε < γ < x. We have V_{ε}(x) = V_{γ}(x) = V(x). So (X, (T_{ε})_{ε∈ℝ^{+}} ) defines a non-Archimedean approach space.

3.10. Non-Archimedean gauges. Finally, we look at the gauges. To a distance operator δ of an approach space X, we associate its unambiguously defined gauge

G = {d ∈ qMet(X) | ∀A ⊆ X, ∀x ∈ X : inf_{a∈A} d(x, a) ≤ δ(x, A)},

where qMet(X) is the collection of all extended quasi-metrics on X. Moreover the distance can be recovered from the gauge by the formula

δ(x, A) = sup_{d∈G} inf_{a∈A} d(x, a).

When the approach space X is given by its tower of closure operators (𝔱_{ε})_{ε∈ℝ^{+}} the gauge can also be described as

G = {d ∈ qMet(X) | ∀ε ≥ 0 : 𝔱_{ε} ≤ 𝔱_{ε}^{d}},

where (𝔱_{ε}^{d})_{ε≥0} is the tower of the approach space (X, δ_{d}).

We recall a few definitions from approach theory [Lowen, 2015]. Given a collection H ⊆ qMet(X) and a quasi-metric d ∈ qMet(X) we say that d is locally dominated by H if for all x ∈ X, ε > 0 and ω < ∞ there exists a d_{x}^{ε,ω} ∈ H such that

d(x, ·) ∧ ω ≤ d_{x}^{ε,ω}(x, ·) + ϵ.  (1)

A subset H of qMet(X) is called locally directed if for any H_{0} ⊆ H finite, we have that sup_{d∈H_{0}} d is locally dominated by H. Given a subset H ⊆ qMet(X) we define

Ĥ := {d ∈ qMet(X) | D locally dominates d}.

Given an approach space X and G its gauge, H ⊆ qMet(X) is called a basis for the gauge G if Ĥ = G. If H is locally directed, then Ĥ is a gauge of an approach space with H as a basis.

3.11. Theorem. Consider a non-Archimedean approach space X with tower of topologies (T_{ε})_{ε∈ℝ^{+}} . Then the associated gauge has a basis consisting of quasi-ultrametrics.

Proof. For n ∈ ℕ_{0} we consider \{\frac{k}{n} | k = 0, ⋯, k = n^{2}\} on [0, n]. At level \frac{k}{n} we choose a finite T_{\frac{k}{n}} -open cover C_{\frac{k}{n}} in such a way that for k > 0 the finite T_{\frac{k−1}{n}} -open cover C_{\frac{k−1}{n}} is a refinement of C_{\frac{k}{n}}. The following notations are frequently used in the setting of quasi-uniform spaces. For x ∈ X and k = 0, ⋯, k = n^{2}

A_{C_{\frac{k}{n}}}^{x} = ⋂\{C | C ∈ C_{\frac{k}{n}} , x ∈ C\}

and

U_{C_{\frac{k}{n}}} = ⋃_{x∈X}\{x\} × A_{C_{\frac{k}{n}}}^{x} .

We employ a standard technique as used with developments in [Lowen, 2015] to construct a function depending on n and on the choice C_{\frac{1}{n}} , ⋯, C_{\frac{n^{2}}{n}} by letting

p_{n} = inf_{k=1}^{n^{2}}(\frac{k − 1}{n} + θ_{U_{C \frac{k}{n}}}) ∧ n,  (2)

where for Z ⊆ X, we use the notation

θ_{Z} : X → [0, ∞] : x ↦ \begin{cases}0 & x ∈ Z \\ ∞ & x ∉ Z. \end{cases}

Clearly for k ∈ \{1, ⋯, n^{2}\} we have U_{C_{\frac{k−1}{n}}} ⊆ U_{C_{\frac{k}{n}}} and every U_{C_{\frac{k}{n}}} is a preorder. This implies that p_{n} is a quasi-ultrametric on X.

Next we show that each p_{n} belongs to the gauge G. Fix ε ≥ 0 and α > ε. Either ε ≥ n and then the open ball B_{p}(x, α) = X ∈ T_{ε}, or ε ∈ [\frac{k−1}{n}, \frac{k}{n}[ for some k ∈ \{1, ⋯, n^{2}\}.

Then for y ∈ A_{C_{\frac{k}{n}}}^{x} we have (x, y) ∈ U_{C_{\frac{k}{n}}} which implies p_{n}(x, y) ≤ \frac{k−1}{n} ≤ ε < α. So

A_{C_{\frac{k}{n}}}^{x} ⊆ B_{p}(x, α) and again B_{p}(x, α) ∈ T_{ε}.

Let H be the collection of all quasi-ultrametrics p_{n}, for arbitrary choices of n and C_{\frac{1}{n}} , ⋯, C_{\frac{n^{2}}{n}} . We prove that H is a basis for the gauge G. Let d ∈ G, x ∈ X and n ∈ ℕ_{0}.

Consider \{\frac{k}{n} | k = 0, ⋯, k = n^{2}\} on [0, n] and at level \frac{k}{n} choose the cover

C_{\frac{k}{n}} = \{B_{d}(x, \frac{k}{n} + \frac{1}{2n}), X\}.

Since d ∈ G the inclusion T d_{\frac{k}{n}} ⊆ T_{\frac{k}{n}} holds, so the cover C_{\frac{k}{n}} is T_{\frac{k}{n}} -open for every k. Moreover clearly at each level C_{\frac{k−1}{n}} refines C_{\frac{k}{n}}. Let p_{n} be the associated quasi-ultrametric as in (2). We show that

d(x, ·) ∧ n ≤ p_{n}(x, ·) + \frac{2}{n}.  (3)

Let y ∈ X. Either p_{n}(x, y) + \frac{2}{n} ≥ n and then we are done, or p_{n}(x, y) + \frac{2}{n} = α ∈ [\frac{k}{n}, \frac{k+1}{n}[ for some k. In this case p_{n}(x, y) = α − \frac{2}{n} ∈ [\frac{k−2}{n}, \frac{k−1}{n}[ which implies (x, y) ∈ U_{C_{\frac{k−1}{n}}} . So

we have y ∈ B_{d}(x, \frac{k−1}{n} + \frac{1}{2n}), by which d(x, y) < \frac{k−1}{n} + \frac{1}{2n} < \frac{k}{n} ≤ α. Since (3) holds for every n ∈ ℕ_{0} it now follows that also (1) is fulfilled for every ε > 0 and ω < ∞, so we can conclude that Ĥ = G.

3.12. Theorem. Consider an approach space X with gauge G, having a basis H consisting of quasi-ultrametrics. Then the associated distance is a non-Archimedean distance.

Proof. The distance δ associated with the gauge G can be derived directly from the basis H by

δ(x, A) = sup_{d∈H} inf_{a∈A} d(x, a).

We only have to show that this distance satisfies (D4_{∨}). Take x ∈ X, A ⊆ X and ε ∈ P_{∨} arbitrary. Then, for any b ∈ A^{(ε)}, d ∈ H and θ > 0, there exists a_{d} ∈ A such that d(b, a_{d}) < ε+θ. Consequently, d(x, a_{d}) ≤ d(x, b)∨d(b, a_{d}) ≤ d(x, b)∨(ε+θ), which proves that inf_{a∈A} d(x, a) ≤ inf_{b∈A^{(ε)}} d(x, b) ∨ (ε + θ). Since this holds for all d ∈ H, it follows that δ(x, A) ≤ δ(x, A^{(ε)}) ∨ ε.

A gauge with a basis consisting of quasi-ultrametrics will be called a non-Archimedean gauge.

Based on the characterisation of non-Archimedean approach spaces in terms of non-Archimedean gauges, we can come back to the embedding of NA-App in App corresponding to the change of base functor C_{φ} : (β, P_{∨})-Cat ⟶ (β, P_{+})-Cat.

3.13. Theorem. NA-App is a concretely reflective subcategory of App. If X is an approach space with gauge G, then its NA-App-reflection 1_{X} : X → X^{u} is given by the approach space X^{u} having G ∩ qMet^{u}(X) as basis for its gauge.

Proof. Since G ∩ qMet^{u}(X) is stable under finite suprema, it is locally directed and therefore Ĝ∩qMet^{u}(X) defines a gauge. Let X^{u} be the associated approach space. Suppose f : X → Y is a contraction with Y in NA-App with a basis H consisting of quasiultrametrics. Then d ∈ H clearly implies d∘f ×f ∈ G∩qMet^{u}(X). By the characterization of a contraction in terms of a gauge basis, we have that f : X^{u} → Y is contractive.

4. The embeddings qMet^{u} ↪ App and Top ↪ App.

Restricting the coreflector App → qMet from [Lowen, 2015] to NA-App, a non-Archimedean approach space X with limit operator λ (distance δ) is sent to its underlying quasiultrametric space (X, d_{λ}) ((X, d_{δ})) given by

d_{λ}(x, y) = λ(ẏ)(x) = δ(x, {y}) = d_{δ}(x, y)

for x, y in X. Moreover restricting the embedding qMet ↪ App from [Lowen, 2015] to qMet^{u}, a quasi-ultrametric space (X, d) is mapped to a non-Archimedean approach space X with limit operator defined by

λ_{d}(U)(x) = sup_{U∈U} inf_{u∈U} d(x, u),

for all U ∈ βX and x ∈ X and distance

δ_{d}(x, A) = inf_{a∈A} d(x, a)

for all A ⊆ X and x ∈ X. We can conclude that qMet^{u} is concretely coreflectively embedded in NA-App and the coreflector is the restriction of the well known coreflector App → qMet.

Considering the lax extension \overline{β} to P_{∨}-Rel, the lax extension of the identity monad 𝕀 to P_{∨}-Rel and the associated morphism (I, \overline{I}) → (β, \overline{β}) of lax extensions, the induced algebraic functor

(β, P_{∨})-Cat ⟶ P_{∨}-Cat,

sends a (β, P_{∨})-algebra (X, a) to its underlying P_{∨}-algebra (X, a · e_{X}), as introduced in section III.3.4 of [Hofmann, Seal, Tholen (eds.), 2014]. Using the isomorphisms described in section 2 and in Theorem 2.4, this functor sends a (β, P_{∨})-algebra (X, a), corresponding to a non-Archimedean approach space X with limit operator λ, to its underlying quasiultrametric space (X, d_{λ}^{−}). This functor has a left adjoint P_{∨}-Cat ↪ (β, P_{∨})-Cat, which associates to a P_{∨}-algebra (X, d), the (β, P_{∨})-algebra corresponding to (X, λ_{d^{−}}).

Next consider the lax homomorphism ι : 2 ⟶ P_{∨}, sending ⊤ to 0 and ⊥ to ∞, which is compatible with the lax extensions of the ultrafilter monad to Rel and P_{∨}-Rel. Analogous to the situation for P_{+} [Hofmann, Seal, Tholen (eds.), 2014] the change-of-base functor associated to the lax homomorphism ι constitutes an embedding (β, 2)-Cat ↪ (β, P_{∨})-Cat. Using the isomorphisms described in section 2, Theorem 2.4, and the well known isomorphism (β, 2)-Cat ≅ Top [Barr, 1970] or [Hofmann, Seal, Tholen (eds.), 2014], this gives an embedding of Top in NA-App. In terms of the limit operator or the distance the embedding Top ↪ NA-App associates the limit operator λ_{T} (distance δ_{T}) to a topological space (X, T ) by λ_{T} U(x) = 0 if U converges to x in (X, T ) (δ_{T} (x, A) = 0 if x ∈ \overline{A}) and with values ∞ in all other cases, for U ∈ βX, A ⊆ X, x ∈ X. Later on we will also make use of the embedding of Top described in terms of the tower. All levels of the approach tower (X, (𝔱_{ε})_{ε≥0}) associated to a topological space (X, T ) coincide, so we have T_{ε} = T for all ε ≥ 0. These formulations of the embedding Top ↪ NA-App are the codomain restrictions of the embedding of Top in App as described in [Lowen, 2015].

ι has a right adjoint p : P_{∨} ⟶ 2, where p(0) = ⊤ and p(v) =⊥ otherwise, that is again a quantale homomorphism. p is also compatible with the lax extensions of the ultrafilter monad β to Rel and P_{∨}-Rel and provides the embedding with a right adjoint (β, P_{∨})-Cat ⟶ (β, 2)-Cat. This functor can also be obtained by restricting the coreflector T : App ⟶ Top, as described in [Lowen, 2015] to NA-App and we will continue in using the notation T. This coreflector sends a non-Archimedean approach space X to a topological space T(X) in which an ultrafilter U converges to a point x precisely when λU(x) = 0 or in which a point x is in the closure of a set A precisely when δ(x, A) = 0. In terms of the tower (T_{ε})_{ε≥0} the topological space T(X) is precisely (X, T_{0}).

Now define the map o : P_{∨} ⟶ 2 by o(v) = ⊤ if and only if v < ∞. Analogous to the situation of P_{+} in [Hofmann, Seal, Tholen (eds.), 2014] the map o is a lax homomorphism, however it is not compatible with the ultrafilter lax extensions. Nevertheless, given a (β, P_{∨})-algebra (X, a), one can still consider the pair (X, oa), where oa : βX⟶ X is defined by U oa x precisely when a(U, x) < ∞. This structure satisfies the reflexivity but not the transitivity condition. In other words, (X, oa) is a pseudotopological space. Now we can apply the left adjoint of the full reflective embedding Top ↪ PsTop to (X, oa) to obtain a topological space and thereby a left adjoint (β, P_{∨})-Cat ⟶ (β, 2)-Cat to the embedding (β, 2)-Cat ↪ (β, P_{∨})-Cat. This functor can also be obtained by restricting the reflector App ⟶ Top, as described in [Lowen, 2015] to NA-App, where the Top-reflection of a non-Archimedean approach space (X, δ) is determined by the non-Archimedean distance associated with the topological reflection of the pretopological closure operator cl, defined by cl(A) := {x ∈ A | δ(x, A) < ∞}.

The results in this section can be summarized in the following diagram.

[figure: NA-App ≅ (β, P_{∨})-Cat; App ≅ (β, P_{+})-Cat; r; Top ≅ (β, 2)-Cat; NA-App ≅ (β, P_{∨})-Cat; r & c; qMet^{u} ≅ P_{∨}-Cat; NA-App ≅ (β, P_{∨})-Cat; c; qMet^{u} ≅ P_{∨}-Cat; qMet ≅ P_{+}-Cat; r; qMet ≅ P_{+}-Cat; App ≅ (β, P_{+})-Cat; c]

5. Initially dense objects in NA-App

At this point, we can introduce two new examples of non-Archimedean approach spaces.

5.1. Example. Consider the quasi-ultrametrics d_{P_{∨}} and d_{P_{∨}}^{−} on P_{∨} defined by

d_{P_{∨}} : P_{∨} × P_{∨} → P_{∨} : (x, y) ↦ x−•y,

and

d_{P_{∨}}^{−} : P_{∨} × P_{∨} → P_{∨} : (x, y) ↦ y−•x.

The first quasi-ultrametric space is a special case of a V-structure on V as introduced in [Hofmann, Seal, Tholen (eds.), 2014]. In this section we show that each of the non-Archimedean approach spaces (P_{∨}, δ_{P_{∨}}) of 3.4 and (P_{∨}, d_{P_{∨}}) and (P_{∨}, d_{P_{∨}}^{−}), as introduced above, are initially dense objects in NA-App. Recall that for a source (f_{i} : X → (X_{i}, λ_{i}))_{i∈I} where (X_{i}, λ_{i}) are given approach spaces in terms of their limit operator, the initial lift on X is described by the limit operator

λ_{in}U(x) = sup_{i∈I} λ_{i}f_{i}(U)(f_{i}(x))

for U ∈ βX and x ∈ X.

5.2. Proposition. For any non-Archimedean approach space X with distance δ and for A ⊆ X, the distance functional

δ_{A} : (X, δ) ⟶ (P_{∨}, δ_{P_{∨}}) : x ↦ δ(x, A)

is a contraction.

Proof. Let x ∈ X and B ⊆ X. By application of 3.3 for B ≠ ∅ and A ≠ ∅ , we have

δ_{P_{∨}}(δ_{A}(x), δ_{A}(B)) = sup_{b∈B} δ(b, A)−•δ(x, A)
≤ sup_{b∈B} δ(b, A)−•(δ(x, B) ∨ sup_{b∈B} δ(b, A))
≤ δ(x, B).

5.3. Theorem. (P_{∨}, δ_{P_{∨}}) is initially dense in NA-App. More precisely, for any non-Archimedean approach space X, the source

(δ_{A} : X ⟶ (P_{∨}, δ_{P_{∨}}))_{A∈2^{X}}

is initial.

Proof. If λ_{in} stands for the initial non-Archimedean limit operator on X, then we already know by Proposition 5.2 that λ_{in} ≤ λ. Conversely, take U ∈ βX and x ∈ X. Then

λ_{in}U(x) = sup_{A∈2^{X}} λ_{P_{∨}}(δ_{A}(U))(δ_{A}(x))
= sup_{A∈2^{X}} sup_{U∈U} δ_{P_{∨}}(δ_{A}(x), δ_{A}(U))
≥ sup_{U∈U} δ_{P_{∨}}(δ_{U}(x), δ_{U}(U))
= sup_{U∈U} δ_{P_{∨}}(δ_{U}(x), {0})
= sup_{U∈U} δ(x, U)
= λU(x).

5.4. Theorem. Both (P_{∨}, d_{P_{∨}}) and (P_{∨}, d_{P_{∨}}^{−}) are initially dense objects in NA-App.

Proof. Since we already know that (P_{∨}, δ_{P_{∨}}) is initially dense in NA-App, it suffices to show that we can obtain it via an initial lift of sources of either of the two objects above. First of all, consider the following source

(g_{α} : (P_{∨}, δ_{P_{∨}}) ⟶ (P_{∨}, δ_{d_{P_{∨}}}))_{α∈ℝ^{+}} ,

where g_{α} is defined as follows:

g_{α} : (P_{∨}, δ_{P_{∨}}) ⟶ (P_{∨}, δ_{d_{P_{∨}}}) : x ↦ \{ \begin{matrix}0 & x > α, \\ α & x ≤ α. \end{matrix}

To show that g_{α} is a contraction, for any α ∈ ℝ^{+}, take x ∈ P_{∨} and B ⊆ P_{∨} arbitrary. For B ≠ ∅, we consider two cases. First suppose x ≤ α, then

δ_{d_{P_{∨}}} (g_{α}(x), g_{α}(B)) = inf_{b∈B} g_{α}(x)−•g_{α}(b) = inf_{b∈B} α−•g_{α}(b) = 0.

In case x > α, we have

δ_{d_{P_{∨}}} (g_{α}(x), g_{α}(B)) = inf_{b∈B} g_{α}(x)−•g_{α}(b) = inf_{b∈B} 0−•g_{α}(b).

If there exists b ∈ B such that b > α, then δ_{d_{P_{∨}}} (g_{α}(x), g_{α}(B)) = 0 ≤ δ_{P_{∨}}(x, B). If for all b ∈ B we have that b ≤ α then sup B ≤ α < x, and thus δ_{d_{P_{∨}}} (g_{α}(x), g_{α}(B)) = α < x = δ_{P_{∨}}(x, B).

It remains to show that the source (g_{α})_{α∈ℝ^{+}} is initial. Let λ_{in} stands for the initial limit operator on P_{∨}, then we know that λ_{in} ≤ λ_{P_{∨}}. To prove the other inequality, take U ∈ βP_{∨} and x ∈ P_{∨} arbitrary. If λ_{P_{∨}}U(x) = 0, then the inequality is clear. In case λ_{P_{∨}}U(x) = x there exists A ∈ U such that x > sup A. Take α arbitrary in the interval [sup A, x[. Then

λ_{d_{P_{∨}}} g_{α}(U)(g_{α}(x)) = sup_{U∈U} inf_{y∈U} d_{P_{∨}}(g_{α}(x), g_{α}(y)) ≥ inf_{y∈A} d_{P_{∨}}(g_{α}(x), g_{α}(y)) = 0−•α = α.

Hence λ_{in}U(x) = sup_{α∈ℝ^{+}} λ_{d_{P_{∨}}} g_{α}(U)(g_{α}(x)) ≥ sup_{α∈[sup A,x[} α = x = λ_{P_{∨}}U(x). So we can conclude that the source is initial.

In a similar way, we can prove that the source

(f_{α} : (P_{∨}, δ_{P_{∨}}) ⟶ (P_{∨}, δ_{d_{P_{∨}}^{−}}))_{α∈ℝ^{+}} ,

with

f_{α} : (P_{∨}, δ_{P_{∨}}) ⟶ (P_{∨}, δ_{d_{P_{∨}}^{−}}) : x ↦ \{ \begin{matrix}α & x > α, \\ 0 & x ≤ α, \end{matrix}

is initial as well.

6. Topological properties on (β, P_{∨})-Cat

In this section we explore topological properties in (β, P_{∨})-Cat, following the relational calculus, developed in V.1 and V.2 of [Hofmann, Seal, Tholen (eds.), 2014] for (𝕋, V)- properties. We introduce low separation properties, Hausdorffness, compactness, regularity, normality and extremal disconnectedness as an application to (β, P_{∨})-Cat of the corresponding (𝕋, V)-properties for arbitrary (𝕋, V)-Cat. For each of the properties p we characterise the property (β, P_{∨})-p in the context NA-App. On the other hand we make use of the well known meaning of these properties in the setting of Top ≅ (β, 2)-Cat. For a non-Archimedean approach space X with tower of topologies (T_{ε})_{ε∈ℝ^{+}} we compare the property (β, P_{∨})-p to the properties:

• X has p at level 0: meaning (X, T_{0}) has p in Top

• X strongly has p: meaning (X, T_{ε}) has p in Top for every ε ∈ ℝ^{+}.

• X almost strongly has p: meaning (X, T_{ε}) has p in Top for every ε ∈ ℝ_{0}^{+}.

6.1. Low separation properties and Hausdorffness. We recall the definition of (β, P_{∨})-p for p the Hausdorff, T_{1} and T_{0} properties in (β, P_{∨})-Cat by giving the pointwise interpretation through the isomorphism in 2.4 in terms of the limit operator. For a study of low separation properties in App we refer to [Lowen, Sioen 2003].

6.2. Definition. A non-Archimedean approach space X is

1. (β, P_{∨})-Hausdorff if λU(x) < ∞ & λU(y) < ∞ ⇒ x = y, for every U ∈ βX and all x, y ∈ X.

2. X is (β, P_{∨})-T1 if λẋ(y) < ∞ ⇒ x = y, for all x, y ∈ X

3. X is (β, P_{∨})-T0 if λẋ(y) < ∞ & λẏ(x) < ∞ ⇒ x = y, for all x, y ∈ X.

Remark that (β, P_{∨})-Hausdorff ((β, P_{∨})-T_{1}, (β, P_{∨})-T_{0}) is equivalent to (β, P_{+})-Hausdorff ((β, P_{+})-T_{1}, (β, P_{+})-T_{0}) as described in [Hofmann, Seal, Tholen (eds.), 2014]

6.3. Theorem. For a non-Archimedean approach space X the following properties are equivalent

1. X is strongly Hausdorff (strongly T_{1}, strongly T_{0} respectively).

2. X is almost strongly Hausdorff (almost strongly T_{1}, almost strongly T_{0} respectively).

3. X is (β, P_{∨})-Hausdorff ((β, P_{∨})-T_{1}, (β, P_{∨})-T_{0})

and imply that X is Hausdorff (T_{1}, T_{0} respectively) at level 0.

Proof. (1) ⇔ (2) This is straightforward.

(1) ⇔ (3) is based on

λU(x) < ∞ ⇔ ∃ε ∈ ℝ^{+}, λU(x) ≤ ε ⇔ ∃ε ∈ ℝ^{+}, U → x in T_{ε},

for U ∈ βX and x ∈ X. The proofs of the other cases follow analogously.

That (β, P_{∨})-Hausdorff ((β, P_{∨})-T_{1}, (β, P_{∨})-T_{0}) is not equivalent to the Hausdorff (T_{1}, T_{0} ) property at level 0 follows from example X_{S} in 3.8 with S = {X, ∅}.

6.4. Compactness. The next property we consider is (β, P_{∨})-compactness. We recall the definition of (β, P_{∨})-compactness in (β, P_{∨})-Cat by giving the pointwise interpretation through the isomorphism in 2.4 in terms of the limit operator.

6.5. Definition. A non-Archimedean approach space X is (β, P_{∨})-compact if

inf_{x∈X} λU(x) = 0,

for all U ∈ βX.

This condition coincides with (β, P_{+})-compactness studied in (β, P_{+})-Cat ≅ App [Hofmann, Seal, Tholen (eds.), 2014] and is equivalent to what is called 0-compactness in [Lowen, 2015]. The proofs of the following results are straightforward.

6.6. Theorem. For a non-Archimedean approach space X the following equivalences hold:

1. (β, P_{∨})-compact ⇔ almost strongly compact

2. compact at level 0 ⇔ strongly compact

That (β, P_{∨})-compactness does not imply compactness at level 0 is well known in the setting of App [Lowen, 2015]. That this is neither the case in the setting of NA-App follows from the following example.

6.7. Example. Consider the example in 3.9 on ]0, ∞[ with T the right order topology. Clearly T is not compact, whereas the topologies at strictly positive levels are all compact.

6.8. Proposition. Let X be a non-Archimedean approach space. If X is (β, P_{∨})-compact and (β, P_{∨})-Hausdorff, then it is a compact Hausdorff topological space.

Proof. Suppose that the non-Archimedean approach space X is both (β, P_{∨})-compact and (β, P_{∨})-Hausdorff. Then (X, T_{ε}) is a compact Hausdorff topological space, for every ε > 0. By the coherence condition of the non-Archimedean tower, we have T_{γ} ⊆ T_{ε} for ε ≤ γ and therefore T_{γ} = T_{ε}, for every γ, ε > 0. Moreover, since T_{0} = ⋁_{γ>0} T_{γ}, all levels of the non-Archimedean tower are equal. This implies that X is topological.

6.9. Regularity. Next we investigate the notion of regularity. We recall the definition of (β, P_{∨})-p for p the regularity property in (β, P_{∨})-Cat, by giving the pointwise interpretation through the isomorphism in 2.4 in terms of the limit operator.

6.10. Definition. A non-Archimedean approach space X is (β, P_{∨})-regular if

λU(x) ≤ λm_{X}(𝔛)(x) ∨ sup_{A∈𝔛,B∈U} inf_{W∈A,b∈B} λW(b),

for all 𝔛 ∈ ββX, U ∈ βX and x ∈ X.

A strong version of regularity was introduced in [Brock, Kent, 1998]. In terms of arbitrary filters it states that λF^{(γ)} ≤ λF ∨ γ, for every F ∈ FX and x ∈ X and in that paper it was shown to be equivalent to regularity at each level. In more generality this condition was also considered in [Colebunders, Mynard, Trott, 2014] in the context of contractive extensions.

The following result gives a characterisation of (β, P_{∨})-regular in terms of the level topologies.

6.11. Theorem. For a non-Archimedean approach space X the following are equivalent:

1. X is strongly regular

2. X is almost strongly regular

3. For all U, W ∈ βX and for all γ ≥ 0: W^{(γ)} ⊆ U ⇒ λU ≤ λW ∨ γ.

4. X is (β, P_{∨})-regular Proof. (1) ⇔ (2) is clear from the fact that ⋃_{ε>0} C_{ε} is a closed basis for the topology T_{0}.

(1) ⇔ (3) is essentially known from [Brock, Kent, 1998].

(3) ⇒ (4). Take 𝔛 ∈ ββX, U ∈ βX and x ∈ X. Put γ = λm_{X}(𝔛)(x) and ε = sup_{A∈𝔛,B∈U} inf_{W∈A,b∈B} λW(b). It is sufficient to assume that both γ and ε are finite. Let 0 < ρ < ∞ be arbitrary and consider

S := {(G, y) | λG(y) ≤ ε + ρ} ⊆ βX × X.

Clearly the filterbasis 𝔛×U has a trace on S, so we can choose R ∈ βS refining this trace. For A ∈ m_{X}(𝔛) and U ∈ U there exist R_{1} ∈ R and R_{2} ∈ R such that A ∈ ⋂_{z∈R_{1}} π_{1}z and U = π_{2}R_{2}, with π_{1} and π_{2} the projections restricted to S. For z ∈ R_{1} ∩ R_{2} we have λ(π_{1}z, π_{2}z) ≤ ε + γ and π_{2}z ∈ A^{(ε+γ)} ∩ U. Finally we can conclude that m_{X}(𝔛)^{(ε+γ)} ⊆ U w! which implies

λU(x) ≤ λm_{X}(𝔛)(x) ∨ (ε + ρ) ≤ γ ∨ (ε + ρ).

By arbitrariness of ρ our conclusion follows.

(4) ⇒ (1). We use a technique similar to the one used in the proof of Theorem 9 in [Brock, Kent, 1998]. Let W be an ultrafilter converging to x ∈ X in T_{γ} for γ ≥ 0 and let U ∈ βX, such that W^{(γ)} ⊆ U. We may assume that γ is finite. Let 0 < ρ < ∞ be arbitrary. Consider

S = {(G, y) | λG(y) ≤ γ + ρ} ⊆ βX × X

and {S_{W} | W ∈ W} with S_{W} = {(G, y) ∈ S | W ∈ G} whenever W ∈ W which is a filterbasis on S. Let S_{W} be the filter generated. Using the restrictions π_{1} and π_{2} of the projections to S, we observe the following facts:

i) π_{2}S_{W} ⊆ W^{(γ)}: This follows from the fact that y ∈ W^{(γ)} implies the existence of G ∈ βX with W ∈ G and λG(y) ≤ γ + ρ.

ii) There exists R ∈ βS satisfying S_{W} ⊆ R and π_{2}R = U: Suppose the contrary, i.e. for every R ∈ βS with S_{W} ⊆ R there exists U_{R} and R ∈ R such that U_{R} ∩π_{2}R = ∅. We can select a finite number of these sets with U_{R_{i}} ∩ π_{2}R_{i} = ∅ and such that ⋃_{i} R_{i} ∈ S_{W}. In view of i) there exists an index j such that π_{2}R_{j} ∈ U which is a contradiction.

iii) With 𝔛 = π_{1}R we have m_{X}(𝔛) = W since W ∈ W implies W ∈ ⋂_{G∈π_{1}S_{W}} G.

Combining these results, we now have:

λU(x) ≤ λm_{X}(𝔛)(x) ∨ sup_{A∈𝔛,B∈U} inf_{V∈A,b∈B} λV(b)
= λW(x) ∨ sup_{R∈R,R′∈R} inf_{V∈π_{1}R,b∈π_{2}R′} λV(b)
≤ γ ∨ sup_{R∈R} inf_{V∈π_{1}R,b∈π_{2}R} λV(b)
≤ γ ∨ sup_{R∈R} inf_{z∈R} λπ_{1}(z)(π_{2}(z))
≤ γ ∨ (γ + ρ).

By arbitrariness of ρ > 0, we can conclude that U converges to x in T_{γ}.

Recall that in (β, P_{+})-Cat ≅ App the notion (β, P_{+})-regularity as described in [Hofmann, Seal, Tholen (eds.), 2014] is equivalent to an approach form of regularity considered in [Robeys, 1992] and [Lowen, 2015] meaning λF^{(γ)} ≤ λF + γ, for every F ∈ FX and x ∈ X. Other forms of regularity obtained by describing the objects of App as relational algebras were studied in [Colebunders, Lowen, Van Opdenbosch, 2016]. Clearly for non-Archimedean approach spaces we have

(β, P_{∨})-regular ⇒ (β, P_{+})-regular ⇒ regular at level 0

The following examples based on the construction in 3.8 show that none of the implications is reversible.

6.12. Example. (1) Let X = {0, 1} and S the Sierpinski topology on X with {1} open. The approach space X_{S} is not (β, P_{+})-regular. This can be seen by taking F = ˙1 and γ = 1. For this choice we have F^{(γ)} = {X} and λ{X}(1) = 2 ≰ λ˙1(1) + 1. However T_{0} is discrete and hence regular.

(2) Let X be infinite and S the cofinite topology on X. The approach space X_{S} is not (β, P_{∨})-regular. However it is (β, P_{+})-regular. To see this let 1 ≤ γ < 2. A filter F on X either contains a finite set and then F^{(γ)} = F and λF^{(γ)} ≤ λF + γ. Or F does not contain a finite set. In that case we have F^{(γ)} = {X} and λF^{(γ)} = 2 ≤ λF + γ. For γ < 1 or 2 ≤ γ the condition λF^{(γ)} ≤ λF + γ is clearly also fulfilled.

6.13. Normality. Next we investigate normality. We recall the definition of normality in (β, P_{∨})-Cat, by giving the pointwise interpretation in terms of the limit operator.

6.14. Definition. A non-Archimedean approach space X is (β, P_{∨})-normal if

â(U, A) ∨ â(U, B) ≥ inf{â(A, W) ∨ â(B, W) | W ∈ βX},

for all ultrafilters U, A, B on X, with â(U, A) = inf{u ∈ [0, ∞] | U^{(u)} ⊆ A}.

Turning the ∨ in the formula into + we have (β, P_{+})-normality as studied in [Hofmann, Seal, Tholen (eds.), 2014]. In case X is a topological (approach) space both notions coincide with

\overline{U} ⊆ A & \overline{U} ⊆ B ⇒ ∃W ∈ βX, \overline{A} ⊆ W & \overline{B} ⊆ W

for all ultrafilters U, A, B on X. As is shown in [Hofmann, Seal, Tholen (eds.), 2014] this condition coincides with the usual notion of normality on the topological (approach) space. An ultrametric approach space is (β, P_{∨})-normal, since by symmetry of the metric the ultrafilter W can be taken to be equal to U. Without symmetry we will encounter examples of quasi-ultrametric approach spaces that are (β, P_{∨})-normal and others that are not.

Next we give some useful characterisations of (β, P_{∨})-normality.

6.15. Proposition. For a non-Archimedean approach space X, the following properties are equivalent:

1. X is (β, P_{∨})-normal

2. â(U, A) < v & â(U, B) < v ⇒ ∃W ∈ βX, â(A, W) < v & â(B, W) < v for all U, A, B ultrafilters on X and v > 0

3. A^{(v)} ∩ B^{(v)} = ∅ ⇒ ∀u < v, ∃C ⊆ X, A^{(u)} ∩ C^{(u)} = ∅ & (X \ C)^{(u)} ∩ B^{(u)} = ∅ for all A, B ⊆ X and v > 0

Proof. That (1) and (2) are equivalent is straightforward. To show that (2) implies (3), let that A^{(v)} ∩ B^{(v)} = ∅ with A, B ⊆ X for some v > 0 and let u < v arbitrary. Suppose on the contrary that for all C ⊆ X, A^{(u)} ∩ C^{(u)} ≠ ∅ or (X \ C)^{(u)} ∩ B^{(u)} ≠ ∅. By Lemma V.2.5.1 in [Hofmann, Seal, Tholen (eds.), 2014] there exist ultrafilters U, A, B on X satisfying

∀U ∈ U : A^{(u)} ∩ U^{(u)} ∈ A & B^{(u)} ∩ U^{(u)} ∈ B.

It follows that â(U, A) ≤ u < v and â(U, B) ≤ u < v. By (2) there exists W ∈ βX with â(A, W) < v & â(B, W) < v. Since A^{(v)} ⊆ W, A^{(u)} ∈ A and A^{(u)^{(v)}} ⊆ A^{(v)^{(v)}} = A^{(v)} we have A^{(v)} ∈ W. In the same way we have B^{(v)} ∈ W which contradicts A^{(v)} ∩ B^{(v)} = ∅. Next we show that (3) implies (2). Let U, A, B be ultrafilters on X, and v > 0 with â(U, A) < v & â(U, B) < v. Choose ε, δ satisfying â(U, A) < ε < δ < v & â(U, B) < ε < δ < v. Let A ∈ A and B ∈ B. We claim that for all C ⊆ X :

A^{(ε)} ∩ C^{(ε)} ≠ ∅ or B^{(ε)} ∩ (X \ C)^{(ε)} ≠ ∅.

Indeed, in view of U^{(ε)} ⊆ A and U^{(ε)} ⊆ B, the assertion ∃C ⊆ X with A^{(ε)} ∩ C^{(ε)} = ∅ & B^{(ε)} ∩ (X \ C)^{(ε)} = ∅ would imply C ∉ U and X \ C ∉ U, which is impossible. By (3) we have A^{(δ)} ∩ B^{(δ)} ≠ ∅. So there exists an ultrafilter W on X refining

{A^{(δ)} ∩ B^{(δ)} | A ∈ A, B ∈ B}.

Clearly W satisfies A^{(δ)} ⊆ W and B^{(δ)} ⊆ W. So we can conclude that â(A, W) < v & â(B, W) < v.

Remark that for a given approach space X, condition (2) in 6.15 implies condition (ii) in Theorem V.2.5.2 of [Hofmann, Seal, Tholen (eds.), 2014] and therefore also (iii) which in [Van Olmen 2005] was shown to be equivalent to approach frame normality of the lower regular function frame of X.

6.16. Theorem. For a non-Archimedean approach space X with tower (T_{ε})_{ε≥0}, consider the following properties

1. X is strongly normal 2. X is almost strongly normal

3. X is (β, P_{∨})-normal

4. X is normal at level 0

The following implications hold: (1) ⇒ (4) and (1) ⇒ (2) ⇒ (3).

Proof. (1) ⇒ (4) and (1) ⇒ (2) are straightforward. We show that (2) ⇒ (3). Suppose (X, T_{u}) is normal, for every u > 0 and let â(U, A) < v & â(U, B) < v for some v > 0. Take u such that â(U, A) < u < v & â(U, B) < u < v. Then for the topological space (X, T_{u}) we have U^{(u)} ⊆ A and U^{(u)} ⊆ B. By the normality of (X, T_{u}) there exists W ∈ βX satisfying A^{(u)} ⊆ W and B^{(u)} ⊆ W. It follows that â(A, W) ≤ u < v and â(B, W) ≤ u < v.

There are no other valid implications between the properties considered in the previous theorem. This is shown by the next examples.

6.17. Example. On X =]0, ∞[, we consider an approach space as in 3.9. We make a particular choice for the topology

T = {B^{c} | B ⊆]0, ∞[, bounded } ∪ {∅},

on X which is finer than the right order topology and consider the approach space (X, (T_{ε})_{ε∈ℝ^{+}} ). The topology T_{0} is not normal since there are no non-empty distinct and disjoint open subsets, although disjoint non-empty closed subsets do exist. So X is not strongly normal either. Let ε > 0 and consider the topological space (X, T_{ε}). It is a normal topological space since for x ≤ ε we have x ∈ A^{(ε)} for every non-empty subset A. So (2) and hence (3) from Theorem 6.16 are satisfied.

6.18. Example. Let X =]0, ∞[ and let (T_{ε})_{ε≥0} the tower as defined in 6.17 starting from T = {B^{c} | B ⊆]0, ∞[, bounded }∪{∅}. We define another tower (S_{γ})_{γ≥0} on X as follows.

S_{γ} = \{ \begin{matrix}P(X) whenever 0 ≤ γ < 1 \\ T_{γ−1} whenever 1 ≤ γ. \end{matrix}

Clearly the tower (S_{γ})_{γ≥0} defines a non-Archimedean approach space on X. For the topology at level 0 we have T_{0} = P(X) is normal, but the topological space (X, S_{1}) = (X, T ) is not normal. So (1) and (2) from Theorem 6.16 do not hold. However the approach space is (β, P_{∨})-normal. Let A and B be non-empty subsets with A^{(v)} ∩ B^{(v)} = ∅ for some v > 0 and let u < v. Clearly v ≤ 1. In that case the level topology for u is discrete. It follows that C = B satisfies the condition in (3) 6.15.

6.19. Example. Consider example (2) in Remark V.2.5.3 in [Hofmann, Seal, Tholen (eds.), 2014]. X = ({x, y, z, w}, d) where d is the quasi-ultrametric structure d(x, z) = d(y, z) = d(w, z) = 1, d(w, x) = d(w, y) = 3, d(x′, x′) = 0 for any x′ ∈ X and d(x′, y′) = ∞ elsewhere. The topology T_{0} is discrete and hence normal. However X = ({x, y, z, w}, d) is not (β, P_{∨})-normal. Let U = ż, A = ẋ and B = ẏ. Then â(U, A) = d(x, z) = 1 and â(U, B) = d(y, z) = 1. For W = ẇ we have â(A, W) = d(w, x) = 3 and â(B, W) = d(w, y) = 3 and for all other choices of W we obtain values ∞. So the space X = ({x, y, z, w}, d) does not satisfy (1) or (2) from 6.16 either.

Next we prove that the notions (β, P_{∨})-normal and (β, P_{+})-normal are unrelated. Again this can be shown by looking at finite non-Archimedean approach spaces that are therefore structured by some quasi-ultrametric.

6.20. Example. Let X = {x, y, z} endowed with the quasi-ultrametric d defined by d(y, x) = 2, d(x, z) = 1, d(y, z) = 1 and d(x′, x′) = 0 for all x′ ∈ X and d(x′, y′) = ∞ elsewhere. Clearly the only inequality that has to be checked is d(x, z) + d(y, z) = 2 ≥ d(y, x) which is no longer valid when + is changed into ∨. The space is (β, P_{+})-normal but not (β, P_{∨})-normal.

6.21. Example. Let X = {x, y, z, w} endowed with the quasi-ultrametric d defined by d(x, z) = 1, d(y, z) = 2, d(w, z) = 2, d(w, y) = 2, d(w, x) = 2 and d(x′, x′) = 0 for all x′ ∈ X and d(x′, y′) = ∞ elsewhere. The space is not (β, P_{+})-normal since d(y, z) + d(x, z) = 3 < 4 = d(w, y) + d(w, x). However X is strongly normal, and therefore (β, P_{∨})- normal. To see this observe that the level topologies for 0 ≤ ε < 1 are discrete and hence normal. For levels 1 ≤ ε < 2 all points are isolated except x for which the smallest neighborhood is {x, z}. This topology is normal since all closed sets are open. For levels 2 ≤ ε we have smallest neighborhoods V_{w} = X, V_{y} = {y, z}, V_{z} = {z} and V_{x} = {x, z}. These levels are normal too since there are no disjoint non-empty closed sets.

6.22. Extremal disconnectedness. Next we investigate extremal disconnectedness. We recall the definition of extremal disconnectedness in (β, P_{∨})-Cat, by giving the pointwise interpretation in terms of the limit operator.

6.23. Definition. A non-Archimedean approach space X is (β, P_{∨})-extremally disconnected if

â(A, U) ∨ â(B, U) ≥ inf{â(W, A) ∨ â(W, B) | W ∈ βX},

for all ultrafilters U, A, B on X, with â(U, A) = inf{u ∈ [0, ∞] | U^{(u)} ⊆ A}.

Turning the ∨ in the formula into + we have (β, P_{+})-extremal disconnectedness as studied in [Hofmann, Seal, Tholen (eds.), 2014]. In case X is a topological (approach) space both notions coincide with

\overline{A} ⊆ U & \overline{B} ⊆ U ⇒ ∃W ∈ βX, \overline{W} ⊆ A & \overline{W} ⊆ B

for all ultrafilters U, A, B on X. As is shown in [Hofmann, Seal, Tholen (eds.), 2014] this condition coincides with the usual notion of extremal disconnectedness on the topological (approach) space. An ultrametric approach space is (β, P_{∨})-extremally disconnected, since by symmetry of the metric the ultrafilter W can be taken to be equal to U. Without symmetry we will encounter examples of quasi-ultrametric approach spaces that are (β, P_{∨})-extremally disconnected and others that are not.

6.24. Proposition. For a non-Archimedean approach space X with tower (T_{ε})_{ε≥0}, the following properties are equivalent:

1. X is (β, P_{∨})-extremally disconnected

2. â(A, U) < v & â(B, U) < v ⇒ ∃W ∈ βX, â(W, A) < v & â(W, B) < v for all U, A, B ultrafilters on X and v > 0

3. A ∩ B = ∅ with A, B ∈ T_{v} ⇒ ∀u < v, A^{(u)} ∩ B^{(u)} = ∅ for all A, B ⊆ X and v > 0.

Proof. The equivalence of (1) and (2) is straightforward. Next we prove that (2) implies (3). If (3) does not hold then there exists v > 0 and A, B ∈ T_{v} with A ∩ B = ∅ and u < v such that A^{(u)} ∩ B^{(u)} ≠ ∅.Take an ultrafilter U containing both A^{(u)} and B^{(u)}. Let U_{T_{u}} be the filter generated by the T_{u}-open sets in U. Then clearly U_{T_{u}} ∨ {A} is a proper filter.

We claim that there exists an A ∈ βX, U_{T_{u}} ∨ {A} ⊆ A with A^{(u)} ⊆ U. Suppose on the contrary that for every such A_{i} ∈ βX with U_{T_{u}} ∨ {A} ⊆ A_{i} there exists A′_{i} ∈ A_{i} with A_{i}^{′(u)} ∉ U. A finite subcollection of ultrafilters {A_{j} | j ∈ J} exists, for which the corresponding sets A′_{j} ∈ A_{j} satisfy ∪_{j∈J}A′_{j} ∈ U_{T_{u}} ∨ {A}. Let U ∈ U_{T_{u}} such that U ∩ A^{(u)} ⊆ ∪_{j∈J}A′_{j}. From A^{(u)} ∈ U we deduce that (U ∩ A)^{(u)} and hence also ∪_{j∈J}A′_{j} belongs to U. A contradiction follows. In the same way there exists an B ∈ βX, U_{T_{u}} ∨ {B} ⊆ B with B^{(u)} ⊆ U. By (2) there exists W ∈ βX, W^{(v)} ⊆ A and W^{(v)} ⊆ B. For every W ∈ W we have W^{(v)} ∩ A ∈ A and W^{(v)} ∩ B ∈ B. By Lemma V.2.5.1 in [Hofmann, Seal, Tholen (eds.), 2014] we get A ∩ C^{(v)} = A ∩ C ≠ ∅ or B ∩ (X \ C)^{(v)} = B ∩ (X \ C) ≠ ∅ which is impossible. Hence we have (3).

Next we prove (3) implies (2). Let A, B, U ∈ βX with â(A, U) < γ & â(B, U) < γ and choose u, v > 0 such that â(A, U) < u < v < γ & â(B, U) < u < v < γ. For A ∈ A ∩ T_{v} and B ∈ B ∩ T_{v} we have A^{(u)} ∩ B^{(u)} ≠ ∅. By (3) we have A ∩ B ≠ ∅. So

{A ∩ B | A ∈ A ∩ T_{v}, B ∈ B ∩ T_{v}}

is a filter basis and a finer ultraflter exists. Applying the second equivalence used in the proof of (ii) ⇔ (iii) in V.2.4.4 of [Hofmann, Seal, Tholen (eds.), 2014] to the topology T_{v} we obtain a W ∈ βX satisfying W^{(v)} ⊆ A and W^{(v)} ⊆ B. So we can conclude that â(W, A) ≤ v < γ & â(W, B) ≤ v < γ.

6.25. Theorem. For a non-Archimedean approach space X with tower (T_{ε})_{ε≥0}, consider the following properties

1. X is strongly extremally disconnected

2. X is almost strongly extremally disconnected

3. X is (β, P_{∨})-extremally disconnected

4. X is extremally disconnected at level 0

The following implications hold: (1) ⇒ (2) ⇒ (3) and (1) ⇒ (4).

Proof. (1) ⇒ (4) and (1) ⇒ (2) are straightforward. We show that (2) ⇒ (3) using characterisation (3) in 6.24. Suppose A ∩ B = ∅ with A, B ∈ T_{v} and let u < v. since (X, T_{v}) is extremally disconnected we have A^{(v)}∩B^{(v)} = ∅ and since T_{v} ⊆ T_{u} the conclusion follows.

There are no other valid implications between the properties considered in the previous theorem. This is shown by the next examples.

6.26. Example. We use the ultrametric space X = ([0, ∞], d_{M}) that played a crucial role in the development of the Banaschewski compactification in [Colebunders, Sioen, 2017]. On [0, ∞] let d_{M} be defined by

d_{M}(x, y) := \begin{cases}x ∨ y & x ≠ y, \\ 0 & x = y. \end{cases}

X clearly is an ultrametric space. When it is considered as an approach space it has a tower (R_{ε})_{ε} of zero dimensional topologies. Since d_{M} is symmetric the space ([0, ∞], d_{M}) is (β, P_{∨})-extremally disconnected. The topology T_{0} is discrete in all points x ≠ 0 and has the usual neighborhood filter in 0. The set A = {1/n | n ≥ 1} is open in T X but its closure is A ∪ {0} which clearly is not open. So T_{0} is not extremally disconnected and hence X is not strongly extremally disconnected. Let ε > 0 and consider the level topology ([0, ∞], R_{ε}). Consider the open set A = {ε + 1/n | n ≥ 1}, then its closure A^{(ε)} = A ∪ [0, ε] which is not open at level ε. So although (3) is fulfilled, none of the other conditions in 6.25 hold.

6.27. Example. Let X = [0, ∞[, endowed with the topology T with neighborhood filters (V(x))_{x∈X} where V(0) is the usual neighborhood filter in the Euclidean topology, and V(x) = ẋ whenever x ≠ 0, so at level 0 we use the same topology as in the previous example 6.26. We define (T_{ε})_{ε∈ℝ^{+}} with T_{0} = T and T_{ε} at level 0 < ε having a neighborhood filter

V_{ε}(x) = \{ \begin{matrix}stack\{[0, x]\} whenever 0 < x ≤ ε \\ V(x) whenever ε < x or x = 0 \end{matrix}

at x ∈ X. The topology at level 0 is ([0, ∞[, T ). As we know from 6.26 it is not extremally disconnected.

Next we consider level ε > 0 and ([0, ∞[, T_{ε}). Let A ∈ T_{ε} then either there is x ≤ ε with x ∈ A. In this case A^{(ε)} = [0, ε] ∪ (A∩]ε, ∞[) which is open, or there is no x ≤ ε with x ∈ A. In this case we have A^{(ε)} = A open. X is extremally disconnected at every strictly positive level, so (2) and (3) from 6.25 are satisfied whereas (1) and (4) are not.

6.28. Example. Let (X, d) be as is 6.19 and now consider (X, d^{−}). Then we have a space with T_{0} extremally disconnected which is not (β, P_{∨})-extremally disconnected, so it satisfies (4), but none of the other conditions in 6.25.

To see that the notions (β, P_{∨})-extremally disconnected and (β, P_{+})-extremally disconnected are unrelated we can consider examples (X, d^{−}) for each of the spaces described in 6.20 and 6.21.

6.29. Topological properties of the initially dense objects. We come back to the examples in 3.4 and 5.1 and investigate their topological properties.

6.30. Example. First we consider X = (P_{∨}, δ_{P_{∨}}). Clearly the limit operator satisfies λU(0) = 0, for any U ∈ βP_{∨}. Thus the space is compact at level 0. P_{∨} is T_{0} since λẋ(y) = 0 = λẏ(x) implies x = y. However, the space is not (β, P_{∨})-T_{0}. To see this, take x, y ∈ [0, ∞[ such that x < y. Then λẋ(y) = y and λẏ(x) = 0. Hence both λẋ(y) and λẏ(x) are finite, but x ≠ y. Since δ_{P_{∨}}(0, A) = 0 for every nonempty subset A, we have 0 ∈ cl A in the topology T_{0}. So clearly T_{0} is neither T_{1} nor regular. Therefore (P_{∨}, δ_{P_{∨}}) is not (β, P_{∨})-T_{1} nor (β, P_{∨})-regular. P_{∨} is strongly normal and strongly extremally disconnected. This follows from the fact that at level 0 (and hence at all other levels) all non-empty closed sets contain 0. We conclude that all levels are normal. On the other hand there are no non-empty disjoint open sets at level 0 since A ∩ A′ = ∅ and ∞ ∈ P_{∨} \ A implies P_{∨} ⊆ P_{∨} \ A. So all levels are extremally disconnected.

6.31. Example. Next we consider the examples (P_{∨}, d_{P_{∨}}) and (P_{∨}, d_{P_{∨}}^{−}) as described in 5.1. For X = (P_{∨}, d_{P_{∨}}) and for any U ∈ βP_{∨}, we have that λ_{d_{P_{∨}}} U(∞) = inf_{U∈U} sup_{y∈U} d_{P_{∨}}(∞, y) = 0, so we get that (P_{∨}, d_{P_{∨}}) is compact at level 0. For x, y ∈ P_{∨} arbitrary, we have that λ_{d_{P_{∨}}} ẋ(y) = 0 if and only if x ≤ y. This proves that (P_{∨}, d_{P_{∨}}) is T_{0}. However (P_{∨}, d_{P_{∨}}) is not (β, P_{∨})-T_{0}. To see this, take x, y ∈ [0, ∞[ such that x < y. Then λ_{d_{P_{∨}}} ẋ(y) = 0, and λ_{d_{P_{∨}}} ẏ(x) = y < ∞. Since d_{P_{∨}}(∞, z) = 0 for every z ∈ P_{∨} we have ∞ ∈ cl A in the topology T_{0} for every nonempty subset A. So clearly T_{0} is not T_{1}. Therefore (P_{∨}, δ_{P_{∨}}) is not (β, P_{∨})-T_{1}. X = (P_{∨}, d_{P_{∨}}) is not (β, P_{+})-regular, since the quasi-ultrametric is not symmetric. Moreover as ∞ ∈ cl A for every subset A, the topology T_{0} is not regular either. X = (P_{∨}, d_{P_{∨}}) is strongly normal and strongly extremally disconnected. To see this, take A ⊆ P_{∨} non-empty. We notice that at level 0 (and hence at all other levels) ∞ ∈ cl(A). So all non-empty closed sets contain ∞. We conclude that all levels are normal. On the other hand all non-empty open sets contain 0. So all levels are extremally disconnected. The results for (P_{∨}, d_{P_{∨}}^{−}) follow analogously.

7. Compact Hausdorff non-Archimedean approach spaces

In the previous section in 6.8 we proved that a non-Archimedean approach space X that is at the same time (β, P_{∨})-compact and (β, P_{∨})-Hausdorff is topological. In this section for a space X we consider the axioms compact at level 0 in combination with Hausdorff at level 0, in the sense that the topology T_{0} is both compact and Hausdorff. In [Lowen, 2015] these conditions are called compact and Hausdorff and in sections 7 and 8 we will use this simplified terminology. We consider NA-App_{2}, the full subcategory of NA-App consisting of all Hausdorff non-Archimedean approach spaces and NA-App_{c2} the full subcategory of NA-App_{2} consisting of all compact Hausdorff non-Archimedean approach spaces. For terminology used in sections 7 and 8 we refer to [Adámek, Herrlich, Strecker, 1990].

NA-App_{2} is epireflective in NA-App and is closed under the construction of finer structures, hence it is a quotient reflective subcategory of NA-App. So NA-App_{2} is monotopological, i.e. for any family (X_{i})_{i∈I} of Hausdorff non-Archimedean approach spaces and any point-separating source (f_{i} : X [figure] X_{i})_{i∈I} there exists a unique initial lift on X. Our next aim in this section is to determine the epimorphisms and extremal monomorphisms in this category.

7.1. Theorem. A contraction f : X ⟶ Y in NA-App_{2} is an epimorphism in NA-App_{2}, if and only if f(X) is T Y -dense.

Proof. One implication is straightforward. If f : X ⟶ Y is a contraction in NA-App_{2} and f(X) is T Y -dense then f : T X ⟶ T Y is an epimorphism in Haus. This implies f is an epimorphism in NA-App_{2} since for any two contractions u, v : Y ⟶ Z in NA-App_{2} with u · f = v · f, also T u · T f = T v · T f in Haus. In order to prove the converse, suppose f : T X ⟶ T Y is a contraction in NA-App_{2} with f(X) not T Y -dense. Set M = cl_{T Y} (f(X)). We use a technique based on the amalgamation, for which we refer to [Dikranjan, Giuli, Tozzi, 1988]. Assume Y has gauge basis H consisting of quasi-ultrametrics. The gauge basis of the amalgamation Y ∐_{M} Y is given by

H_{Y ∐_{M} Y} = {u_{g_{d}} | d ∈ H} ↓,

where

u_{g_{d}}(\overline{x_{i}}, \overline{y_{j}}) = \{ \begin{matrix}inf_{m∈M} d(x, m) ∨ d(m, y) & i ≠ j; x, y ∉ M, \\ d(x, y) & elsewhere. \end{matrix}

Moreover since M is cl_{T Y} -closed in Y, the space Y ∐_{M} Y is Hausdorff. Both statements follow analogously to the construction and results in [Claes, Colebunders, Gerlo, 2007]. Now, consider the diagram

X ⟶^{f} Y [figure: j_{1}; j_{2}] Y ∐ Y ⟶^{φ} Y ∐ _{M}Y.

Let u = φ · j_{1} and v = φ · j_{2}. Then u · f = v · f, but u ≠ v. Hence f is not an epimorphism in NA-App_{2}.

We proceed by determining the extremal monomorphisms in NA-App_{2}. Since cl_{T} is idempotent and weakly hereditary, NA-App_{2} is an (E^{cl^{T}}, M^{cl^{T}})-category, where E^{cl^{T}} = Epi(NA-App_{2}). The following theorem determines the extremal monomorphisms in NA-App_{2}.

7.2. Theorem. The following classes of morphisms in NA-App_{2} coincide:

1. The class of all regular monomorphisms;

2. The class of all extremal monomorphisms;

3. The class of all cl_{T}-closed embeddings.

Proof. To prove that extremal monomorphisms are cl_{T}-closed embeddings, take E^{cl^{T}} the class of all cl_{T}-dense contractions and M^{cl^{T}} the class of all cl_{T}-closed embeddings. By the Theorem in section 2.4 in [Dikranjan, Tholen, 1995], we have that NA-App_{2} is an (E^{cl^{T}}, M^{cl^{T}})-category. Since by 7.1 we have E^{cl^{T}} = Epi(NA-App_{2}), it follows that ExtrMono(NA-App_{2}) ⊆ M^{cl^{T}} [Adámek, Herrlich, Strecker, 1990].

In order to prove that every cl_{T}-closed embedding is a regular monomorphism, take f : X ⟶ Y a cl_{T Y} -closed embedding in NA-App_{2}. Consider the construction of the amalgamation of Y with respect to f(X).

X ⟶^{f} Y [figure: j_{1}; j_{2}] Y ∐ Y ⟶^{φ} Y ∐ _{f(X)}Y.

Since f(X) is cl_{T Y} -closed, the amalgamation is Hausdorff. f is the equalizer of the pair (φ · j_{1}, φ · j_{2}). Hence f is a regular monomorphism.

We recall the following definitions from section 8.1 in [Dikranjan, Tholen, 1995]. Consider a functor F : X ⟶ Y. For an object Y ∈ Y, let F^{−1}Y := {X ∈ X | F X = Y } be the class of objects of the fibre of F at Y, and let F̃^{−1}Y := {X ∈ X | F X ≅ Y } be its replete closure in Y. F is called transportable if for every B ∈ F̃^{−1}Y, there is A ∈ F^{−1}Y with A ≅ B.

7.3. Proposition. The restriction of the coreflector T : NA-App_{2} ⟶ Haus is transportable.

Proof. Take (Y, T ) an arbitrary object in Haus. Take (X, (S_{ε})_{ε∈ℝ^{+}} ) ∈ T̃^{−1}(Y, T ) arbitrary. There exists an isomorphism f : (X, S_{0}) ⟶ (Y, T ) in Haus. For any ε > 0, define the topology on Y by transportation of the topology on X by putting T_{ε} = {f(A) | A ∈ S_{ε}} and let T_{0} = T . Then (T_{ε})_{ε∈ℝ^{+}} is a non-Archimedean tower on Y in the fibre of (Y, T ) and f : (X, (S_{ε})_{ε∈ℝ^{+}} ) ⟶ (Y, T_{ε})_{ε∈ℝ^{+}} ) is an isomorphism in NA-App_{2}.

Now we can conclude the following.

7.4. Theorem. NA-App_{2} is cowellpowered.

Proof. By Theorem 7.1 we have T ( Epi(NA-App_{2})) ⊆ Epi(Haus). Since T is fibre small and transportable 7.3, the Theorem from section 8.1 in [Dikranjan, Tholen, 1995] implies cowellpoweredness of NA-App_{2}.

7.5. Theorem. NA-App_{c2} is an epireflective subcategory of NA-App_{2}.

Proof. By 3.13 NA-App is closed in App under the formation of products and subspaces. Since the class of compact Hausdorff approach spaces is known to be closed under the formation of products and closed subspaces in App [Lowen, 2015] we can conclude that NA-App_{c2} is closed under formation of products and closed subspaces in NA-App_{2}. Since by 7.2 and 7.4 NA-App_{2} is a cowellpowered (Epi, Extremal mono)-category, NA-App_{c2} is epireflective in NA-App_{2}.

8. Non-Archimedean Hausdorff Compactifications

Due to the foregoing theorem 7.5, there is a categorical construction of an epireflector E : NA-App_{2} → NA-App_{c2} with epireflection morphisms e_{X} : X ⟶ KX, for every X in NA-App_{2}. The question remains whether the epireflection morphisms e_{X} are embeddings. The following proposition shows that in general, this is not the case.

8.1. Proposition. A Hausdorff non-Archimedean approach space X that can be embedded in a compact Hausdorff non-Archimedean approach space Y has a topological coreflection T X that is a Tychonoff space.

Proof. If f : X ⟶ Y is an embedding with Y ∈ NA-App_{c2}, then T f : T X ⟶ T Y is an embedding in Top with T Y compact Hausdorff topological. So T X is a Tychonoff space.

In theorem 8.3 below we will formulate sufficient conditions on a Hausdorff non-Archimedean approach space X, to ensure that there exists an embedding into a compact Hausdorff non-Archimedean approach space. Our construction is based on Shanin’s compactification of topological spaces.

Given 𝔖, a collection of closed sets of a topological space X satisfying the conditions (i) ∅, X ∈ 𝔖 and (ii) G_{1}, G_{2} ∈ 𝔖 ⇒ G_{1} ∪ G_{2} ∈ 𝔖, a 𝔖-family is a non-empty collection of sets belonging to 𝔖 satisfying the finite intersection property (f.i.p.). A 𝔖-family F is called vanishing if ⋂_{F ∈F} F = ∅. A maximal 𝔖-family is a 𝔖-family which is not contained in any 𝔖-family as a proper subcollection. If 𝔖 moreover is a closed basis of X, then a compact topological space σ(X, 𝔖) = (S, S) in which X is densely embedded is constructed in [Nagata, 1968] on the set S = X ∪ X′ with X′ the set of all maximal vanishing 𝔖-families. S is endowed with the topology S with {S(G) | G ∈ 𝔖} as a closed basis, where S(G) = G ∪ {p ∈ X′ | G ∈ p}.

Given a non-Archimedean approach space X we will construct an extension of X by first considering a special closed base 𝔖 for T X and constructing σ(T X, 𝔖).

8.2. Theorem. Any non-Archimedean approach space X, given by its tower of topological spaces (T_{ε})_{ε∈ℝ^{+}} and tower of closed sets (C_{ε})_{ε∈ℝ^{+}} , can be densely embedded in a compact non-Archimedean approach space Σ(X, 𝔖), constructed from the closed basis 𝔖 = ⋃_{ε>0} C_{ε} of T X and such that the topological coreflection T Σ(X, 𝔖) is the Shanin compactification σ(T X, 𝔖) of the topological coreflection T X.

Proof. Let X be a non-Archimedean approach space given by its tower of topological spaces (T_{ε})_{ε∈ℝ^{+}} and tower of closed sets (C_{ε})_{ε∈ℝ^{+}} , and take 𝔖 = ⋃_{ε>0} C_{ε}. By the coherence condition of the non-Archimedean tower, 𝔖 is a closed basis for the topological space T X = (X, T_{0}) and moreover 𝔖 clearly satisfies the assumptions (i) and (ii) made above. Let σ(T X, 𝔖) = (S, S) be Shanin’s compactification.

For every ε ∈ ℝ^{+} the collection {S(G) | G ∈ C_{ε}} is a basis for the closed sets D_{ε} of a topology R_{ε} on S. However for (R_{ε})_{ε∈ℝ^{+}} the coherence condition from 3.7 need not be satisfied. In order to force the coherence condition we define the following tower of topologies on S; for α ∈ ℝ^{+}, let

S_{α} = ⋁_{β>α} R_{β},

with the supremum taken in Top. The set S endowed with the tower (S_{ε})_{ε∈ℝ^{+}} defines a non-Archimedean approach space which we denote by Σ(X, 𝔖).

Next we investigate its topological coreflection T Σ(X, 𝔖) = (S, S_{0}). By definition, the collection ⋃_{β>0} D_{β} is a closed basis for (S, S_{0}) and therefore the collection

{S(G) | G ∈ C_{ε}, for some ε > 0} = {S(G) | G ∈ 𝔖}

is a basis for its closed sets too. It follows that (S, S_{0}) = σ(T X, 𝔖), so Σ(X, 𝔖) is a compact non-Archimedean approach space.

Let j : X → S be the canonical injection. By construction, for ε ∈ ℝ^{+} the map j : (X, T_{ε}) → (S, R_{ε}) is an embedding in Top. Since T_{ε} = ⋁_{γ>ε} T_{γ} the source (j : (X, T_{ε}) → (S, R_{γ}))_{γ>ε} is initial in Top which then implies that j : (X, T_{ε}) → (S, S_{ε}) is initial too. So finally we have that j : (X, (T_{ε})_{ε∈ℝ^{+}} ) ⟶ (Y, (S_{ε})_{ε∈ℝ^{+}} ) is an embedding in NA-App. That the embedding j is dense follows already from the result at level 0.

In general, this compactification is not Hausdorff. To ensure the Hausdorff property, we need stronger conditions on X.

8.3. Theorem. Let X be a non-Archimedean approach space given by its tower of closed sets (C_{ε})_{ε∈ℝ^{+}} . The compactification described in Theorem 8.2 is Hausdorff if and only if the following conditions are fulfilled

i) X is Hausdorff,

ii) ∀ε > 0 : ∀G ∈ C_{ε}, ∀x ∉ G : ∃ 0 < γ ≤ ε, ∃H ∈ C_{γ} such that x ∈ H and H ∩ G = ∅,

iii) ∀ε > 0 : ∀F, G ∈ C_{ε}, F ∩ G = ∅ : ∃ 0 < γ ≤ ε, ∃H, K ∈ C_{γ} such that F ∩ H = ∅, G ∩ K = ∅, H ∪ K = X.

Proof. Remark that the collection 𝔖 = ⋃_{ε>0} C_{ε} used in the construction of 8.2 is closed under finite intersections. Using this fact, condition (ii) is equivalent to condition C) and (iii) is equivalent to condition D) in [Nagata, 1968]. So Theorem IV.3 C) and D) in [Nagata, 1968] can be applied.

8.4. Corollary. For a non-Archimedean approach space X that is Hausdorff, (β, P_{∨})- regular and (β, P_{∨})-normal, the compactification described in Theorem 8.2 is Hausdorff.

Proof. Condition (i) in 8.3 is fulfilled. By 6.11 all strictly positive level topologies are regular, so in (ii) of 8.3 we can take γ = ε. In order to see that also (iii) is fulfilled let ε > 0 and F, G ∈ C_{ε}. Let 0 < γ < ε be arbitrary. Since X is (β, P_{∨})-normal, by 6.15 there exists C ⊆ X with F ∩ C^{(γ)} = ∅ and (X \ C)^{(γ)} ∩ G = ∅. Then H = C^{(γ)} and K = (X \ C)^{(γ)}, satisfy the conditions needed in (iii).

Remark that in the previous corollary the condition (β, P_{∨})-normal can be weakened to Van Olmen’s normality, condition V.2.5.2(iii) in [Hofmann, Seal, Tholen (eds.), 2014]. Note that if X is a topological (non-Archimedean approach) space, T_{ε} = T_{0} and C_{ε} = C_{0}, for every ε > 0. In this case condition iii) is equivalent to X being a normal topological space.

We get the following corollary.

8.5. Corollary. Let X be a non-Archimedean approach space such that the conditions from theorem 8.3 are fulfilled. Then the topological coreflection T X is Tychonoff.

Proof. The conditions from Theorem 8.3 imply that X can be embedded in a compact non-Archimedean approach space Σ(X, 𝔖), with T Σ(X, 𝔖) the Shanin compactification of T X, a compact Hausdorff topological space.

As the Shanin compactification need not be an reflection, we end this section by noting that in general the dense embedding j : X → Σ(X, 𝔖) constructed above, is not a reflection.

References

J. Adámek, H. Herrlich and G.E. Strecker, Abstract and concrete categories. Pure and Applied Mathematics, John Wiley and Sons Inc., New York, 1990.

M. Barr, Relational algebras. In Reports of the Midwest Category Seminar, IV, Lect. Notes Math., Springer, Berlin, 137, (1970), 39-55.

P. Brock and D. Kent, On Convergence Approach Spaces. Appl. Categ. Struct. 6:117 (1998).

V. Claes, E. Colebunders and A. Gerlo, Epimorphisms and cowellpoweredness for separated metrically generated theories. Acta Math. Hungar., 114, (2007), 133-152.

M. M. Clementino and D. Hofmann, Topological features of lax algebras, Appl. Categ. Struct., 11, (2003) 267-286.

E. Colebunders, F. Mynard and W. Trott, Function spaces and contractive extensions in approach theory: the role of regularity, Appl. Categ. Struct., 22, (2014), 551-563.

1484  EVA COLEBUNDERS AND KAREN VAN OPDENBOSCH

E. Colebunders, R. Lowen and K. Van Opdenbosch: Regularity for relational algebras and

approach spaces, Topol. Appl. 200, (2016), 79-100, doi 10.1016/j.topol.2015.12.014. E. Colebunders, M. Sioen: The Banaschewski compactification revisited, preprint, 2017 J. de Groot, Non-Archimedean metrics in topology, Proc. AMS 7,6 (1956) 948-956. D. Dikranjan, E. Giuli and A. Tozzi, Topological categories and closure operators. Quaes-

tiones Math., 11(3) (1988), 323-337. D. Dikranjan and W. Tholen, Categorical structure of closure operators. Mathematics

and its applications 346, Kluwer Academic Publishers Group, Dordrecht, 1995. D. Hofmann, G.J. Seal and W. Tholen (eds.), Monoidal Topology, A categorical Approach

to Order, Metric and Topology. Cambridge University Press, 2014. F.W. Lawvere, Metric spaces, generalized logic, and closed categories, Rendiconti del

Seminario Matematico e Fisico di Milano 43 (1973) 135-166. R. Lowen, Index Analysis: Approach Theory at Work. Springer, Berlin, 2015. R. Lowen, M. Sioen, A Note on Separation in AP, Applied General Topology, 4(2) (2003),

475-486. A.F. Monna, Remarques sur les métriques non-Archimédiennes I,II, Indag. Math. 12,

(1950)122-133, 179-191. J. Nagata, Modern general topology, Bibliotheca Mathematica, Vol. VII, 1968. K. Robeys, Extensions of Products of Metric Spaces. PhD Thesis, Universiteit Antwerpen,

1992. C. Van Olmen, A study of the interaction between Frame theory and approach theory, PhD

thesis, University of Antwerp, 2005

Vakgroep Wiskunde, Vrije Universiteit Brussel Pleinlaan 2, 1050 Brussel

Email: evacoleb@vub.ac.be
      karen.van.opdenbosch@vub.ac.be
This article may be accessed at http://www.tac.mta.ca/tac/

THEORY AND APPLICATIONS OF CATEGORIES (ISSN 1201-561X) will disseminate articles that significantly advance the study of categorical algebra or methods, or that make significant new contributions to mathematical science using categorical methods. The scope of the journal includes: all areas of pure category theory, including higher dimensional categories; applications of category theory to algebra, geometry and topology and other areas of mathematics; applications of category theory to computer science, physics and other mathematical sciences; contributions to scientific knowledge that make use of categorical methods. Articles appearing in the journal have been carefully and critically refereed under the responsibility of members of the Editorial Board. Only papers judged to be both significant and excellent are accepted for publication. Full text of the journal is freely available from the journal’s server at http://www.tac.mta.ca/tac/. It is archived electronically and in printed paper format.

Subscription information Individual subscribers receive abstracts of articles by e-mail as they are published. To subscribe, send e-mail to tac@mta.ca including a full name and postal address. For institutional subscription, send enquiries to the Managing Editor, Robert Rosebrugh, rrosebrugh@mta.ca.

Information for authors The typesetting language of the journal is TEX, and L^{A}TEX2e is required. Articles in PDF format may be submitted by e-mail directly to a Transmitting Editor. Please obtain detailed information on submission format and style files at http://www.tac.mta.ca/tac/.

Managing editor. Robert Rosebrugh, Mount Allison University: rrosebrugh@mta.ca

TEXnical editor. Michael Barr, McGill University: barr@math.mcgill.ca

Assistant TEX editor. Gavin Seal, Ecole Polytechnique Fédérale de Lausanne:

gavin seal@fastmail.fm

Transmitting editors.

Clemens Berger, Université de Nice-Sophia Antipolis: cberger@math.unice.fr Richard Blute, Université d’ Ottawa: rblute@uottawa.ca Lawrence Breen, Université de Paris 13: breen@math.univ-paris13.fr Ronald Brown, University of North Wales: ronnie.profbrown(at)btinternet.com Valeria de Paiva: Nuance Communications Inc: valeria.depaiva@gmail.com Ezra Getzler, Northwestern University: getzler(at)northwestern(dot)edu Kathryn Hess, Ecole Polytechnique Fédérale de Lausanne: kathryn.hess@epfl.ch Martin Hyland, University of Cambridge: M.Hyland@dpmms.cam.ac.uk Anders Kock, University of Aarhus: kock@imf.au.dk Stephen Lack, Macquarie University: steve.lack@mq.edu.au F. William Lawvere, State University of New York at Buffalo: wlawvere@buffalo.edu Tom Leinster, University of Edinburgh: Tom.Leinster@ed.ac.uk Ieke Moerdijk, Utrecht University: i.moerdijk@uu.nl Susan Niefield, Union College: niefiels@union.edu Robert Paré, Dalhousie University: pare@mathstat.dal.ca Jiri Rosicky, Masaryk University: rosicky@math.muni.cz Giuseppe Rosolini, Università di Genova: rosolini@disi.unige.it Alex Simpson, University of Ljubljana: Alex.Simpson@fmf.uni-lj.si James Stasheff, University of North Carolina: jds@math.upenn.edu Ross Street, Macquarie University: ross.street@mq.edu.au Walter Tholen, York University: tholen@mathstat.yorku.ca Myles Tierney, Université du Québec à Montréal : tierney.myles4@gmail.com R. J. Wood, Dalhousie University: rjwood@mathstat.dal.ca